An aqueous solution is 6.00 % methanol CH3OH by mass with a density = 0.988 g/mL. What is the molarity of CH3OH in this solution?


1
Expert's answer
2021-01-04T03:48:30-0500

To calculate molarity we need to first find the number of moles present.

Mass % of methanol CH3OH is 6%

That means for every 100g of solution ther are 6g of CH3OH


Molar mass of CH3OH = 12 + (4x1) + 16 = 32g/mol


Moles of CH3OH in 6g = "\dfrac{mass}{Molar Mass} = \dfrac{6g}{32g/mol} = 0.1875mol" of CH3OH


Now we need to finde the volume of solution


Density = 0.988g/mL


Mass of solution is 100g


Volume = "\dfrac{mass}{density} = \dfrac{100g}{0.988g/mL} = 101.2146mL"


0.187moles are in 101.2146mL

? moles will be in 1000mL of solution


Molarity = "\dfrac{0.187mol x 1000mL}{101.2146mL} = 1.848M" CH3OH